OpenStudy (anonymous):

Medals for whoever helps with all question 7.You earn $245 a week in gross pay, and 2/5 of your check is taken out for taxes and deductions. How much of your check is taken out for taxes and deductions? A.$49 B.$98 C.$150 8.A minor league baseball ticket price decreased $2.50 for each of two years, but then it increased $4.25 in each of the next three years. What was the overall increase of the ticket price? A.$7.75 B.$12.75 C.$3.50

OpenStudy (anonymous):

7.B and 8.A:)

OpenStudy (amistre64):

2(245) ------ = 2(49) = 98 5

OpenStudy (anonymous):

tht was actually rlly easy..my lil sister can do it:)

OpenStudy (amistre64):

i have a trained monkey under the sink that can do it.... so yeah!! lol

OpenStudy (anonymous):

answer (1) is B \[\frac{245}{1} \times \frac{2}{5}=\frac{490}{5} = 98\]answer(2) is A\[2\times-2.50+3\times4.25\]\[-5+12.75=7.75\]